OpenStudy (anonymous):

Which expression(s) is equal to 5^10 Expression 1:(5^-3 x 5^-2)5 Expression 2:((5^-5) / (5^-2 x 5^-1)) 5

OpenStudy (anonymous):

1.) = 5^(-3-2)*5^1=5^(-5+1)

OpenStudy (chelseatheruler):

5ever young

OpenStudy (anonymous):

2.) = 5^(-5+2+1+1)